OpenStudy (txslabrat):

A fireworks company has two types of rockets called Zinger 1 and Zinger 2. The polynomial -16t2 + 150t gives the height in feet of Zinger 1 at t seconds after launch. The polynomial -16t2 + 165t gives the height of Zinger 2 at t seconds after launch. If the rockets are launched at the same time and both explode 6 seconds after launch, how much higher is Zinger 2 than Zinger 1 when they explode?

OpenStudy (anonymous):

-576+900= 324 -576+990=414 414-324=90. thus 90 feet